OpenStudy (anonymous):

Simplify 6 sqrt[44] - sqrt[45] - sqrt[99] + sqrt[20]

OpenStudy (anonymous):

i got -6 sqrt[80]

OpenStudy (anonymous):

\[6\sqrt{44 }- \sqrt{45} -\sqrt{99} + \sqrt{20} \]\[6\left(2 \sqrt{11}\right)- 3 \sqrt{5} -3 \sqrt{11} + 2 \sqrt{5} \]\[6\left(2 \sqrt{11}\right)- 3 \sqrt{5} -3 \sqrt{11} + 2 \sqrt{5} \]\[\left(- 3 \sqrt{5}+ 2 \sqrt{5}\right)+\left(6\left(2 \sqrt{11}\right)-3 \sqrt{11}\right) \]\[-\sqrt{5}+9 \sqrt{11} \]

OpenStudy (anonymous):

Sorry. The second and third lines are duplicates.